# Electromagnetic polarizabilities of the triplet hadrons in heavy hadron chiral perturbation theory

### v1

@@ Abstract
We investigate the electromagnetic polarizabilities of singly heavy mesons and doubly heavy baryons within the framework of heavy hadron chiral perturbation theory up to $\mathcal{O}(p^{3})$. We estimate the low-energy constants using the non-relativistic constituent quark model. A striking prediction of our study is the giant electric polarizabilities of the $D^{*}$ mesons: $\alpha_{E}(\bar{D}^{*0})\approx 291.4\times 10^{-4}\text{ fm}^{3}$ and $\alpha_{E}(D^{*-})\approx-0.4-64.4i\times 10^{-4}\text{ fm}^{3}$. These anomalously large values arise from the near-degenerate mass between $D^{*}$ and $D\pi$, which are orders of magnitude larger than those of their bottom counterparts. This kinematic coincidence induces a pronounced cusp structure in the chiral loops, reflecting the long-range dynamics of a pion cloud. For doubly heavy baryons, polarizabilities depend strongly on heavy-flavor composition: the bcqbcq system differs markedly from ccqccq and bbqbbq due to mixing with scalar heavy-diquark states. Using heavy diquark–antiquark symmetry (HDAS), we unify the chiral dynamics of singly heavy mesons and doubly heavy baryons in the heavy-quark limit. The pion-loop contributions dominate the electromagnetic structure of heavy hadrons and provide essential benchmarks for future lattice QCD simulations.
@@ III.1 NUMERICAL RESULTS
The numerical results for the electromagnetic polarizabilities of singly heavy mesons are shown in Table 3. For pseudoscalar mesons, we observe that $\alpha_{E}^{(c_{1}-g_{1})}(P)=\beta_{M}^{(c_{1}-g_{1})}(P)=0$ due to the absence of $\phi PP$ vertex. The results indicate that the long‑range chiral corrections from the $P^{*}\phi$-loop shown in Fig. 1($c_{2}$-$g_{2}$) provide the dominant contribution to the electric polarizabilities, whereas the magnetic-dipole transition $P^{*}\to P\gamma$ shown in Fig. 1($b_{2}$) mainly determines the magnetic polarizabilities.
Table 3: The numerical results of singly heavy meson electromagnetic polarizabilities (in unit of $10^{-4}~\mathrm{fm}^{3}$).
@@ IV.1 NUMERICAL RESULTS
Table 5: The numerical results of doubly heavy baryon electromagnetic polarizabilities (in unit of $10^{-4}~\mathrm{fm}^{3}$). $b[cq]$ and $b\{cq\}$ denote scalar $S_{[cq]}=1$ state $\mathcal{T}$ and axial-vector $S_{\{cq\}}=1$ state $\mathcal{B}$, respectively.
@@
The numerical results for the electromagnetic polarizabilities of doubly heavy baryons are shown in Table 5. For the baryons containing two identical heavy quarks (ccqccq and bbqbbq), the electric polarizabilities exhibit a pattern where the leading loop contributions are governed by the spin-changing intermediate states. Specifically, the spin-1/2 baryon $\mathcal{B}$ is predominantly influenced by the $\mathcal{B}^{*}\phi$ loops ($\alpha_{E}^{(c_{2}-g_{2})}$), while its spin-3/2 counterpart $\mathcal{B}^{*}$ is mainly affected by the $\mathcal{B}\phi$ loops. This pattern contrasts with observations in singly heavy baryons and light-flavor baryons [69, 22]. The discrepancy arises from differences in the spin–flavor structures of the different baryon multiplets, which result in distinct axial-vector couplings within the chiral loop diagrams.

### v2

@@ Abstract
We investigate the electromagnetic polarizabilities of singly heavy mesons and doubly heavy baryons within the framework of heavy hadron chiral perturbation theory up to $\mathcal{O}(p^{3})$. We estimate the low-energy constants using the non-relativistic constituent quark model. A striking prediction of our study is the giant electric polarizabilities of the $D^{*}$ mesons: $\alpha_{E}(\bar{D}^{*0})\approx 294\times 10^{-4}\text{ fm}^{3}$ and $\alpha_{E}(D^{*-})\approx 1.42-64.5\text{i}\times 10^{-4}\text{ fm}^{3}$. These anomalously large values arise from the near-degenerate mass between $D^{*}$ and $D\pi$, which are orders of magnitude larger than those of their bottom counterparts. This kinematic coincidence induces a pronounced cusp structure in the chiral loops, reflecting the long-range dynamics of a pion cloud. For doubly heavy baryons, polarizabilities depend strongly on heavy-flavor composition: the bcqbcq system differs markedly from ccqccq and bbqbbq due to mixing with scalar heavy-diquark states. Using heavy diquark–antiquark symmetry (HDAS), we unify the chiral dynamics of singly heavy mesons and doubly heavy baryons in the heavy-quark limit. The pion-loop contributions dominate the electromagnetic structure of heavy hadrons and provide essential benchmarks for future lattice QCD simulations.
@@ III.1 NUMERICAL RESULTS
The numerical results for the electromagnetic polarizabilities of singly heavy mesons are shown in Table 3. The uncertainties are derived through standard error propagation. They originate from the experimental uncertainties in the axial coupling constants $g$ and the physical meson masses, as well as the theoretical uncertainty associated with the constituent quark masses. For pseudoscalar mesons, we observe that $\alpha_{E}^{(c_{1}-g_{1})}(P)=\beta_{M}^{(c_{1}-g_{1})}(P)=0$ due to the absence of $\phi PP$ vertex. The results indicate that the long‑range chiral corrections from the $P^{*}\phi$-loop shown in Fig. 1($c_{2}$)-($g_{2}$) provide the dominant contribution to the electric polarizabilities, whereas the magnetic-dipole transition $P^{*}\to P\gamma$ shown in Fig. 1($b_{2}$) mainly determines the magnetic polarizabilities.
Table 3: The numerical results of singly heavy meson electromagnetic polarizabilities (in unit of $10^{-4}~\mathrm{fm}^{3}$). The values in parentheses represent the uncertainties of the results.
@@ IV.1 NUMERICAL RESULTS
Table 5: The numerical results of doubly heavy baryon electromagnetic polarizabilities (in unit of $10^{-4}~\mathrm{fm}^{3}$). $b[cq]$ and $b\{cq\}$ denote scalar $S_{[cq]}=1$ state $\mathcal{T}$ and axial-vector $S_{\{cq\}}=1$ state $\mathcal{B}$, respectively. Uncertainties are given in parentheses.
@@
The error estimates provided in Table 5 account for the uncertainties in the input parameters, specifically the axial coupling constants ($g_{i}$, $c_{i}$) and the constituent quark magnetic moments $\mu_{q}$. The numerical results for the electromagnetic polarizabilities of doubly heavy baryons are shown in Table 5. For the baryons containing two identical heavy quarks (ccqccq and bbqbbq), the electric polarizabilities exhibit a pattern where the leading loop contributions are governed by the spin-changing intermediate states. Specifically, the spin-1/2 baryon $\mathcal{B}$ is predominantly influenced by the $\mathcal{B}^{*}\phi$ loops ($\alpha_{E}^{(c_{2}-g_{2})}$), while its spin-3/2 counterpart $\mathcal{B}^{*}$ is mainly affected by the $\mathcal{B}\phi$ loops. This pattern contrasts with observations in singly heavy baryons and light-flavor baryons [69, 22]. The discrepancy arises from differences in the spin–flavor structures of the different baryon multiplets, which result in distinct axial-vector couplings within the chiral loop diagrams.